Please help.
A rushed subway traveler got on an up-escalator and also walked up. After 45 seconds he realized that he had left his briefcase at the bottom. Since he did not have access to a down-escalator, he turned and walked down the up-escalator reaching the bottom in two minutes. If he had traveled up 288 feet, find the rate of the escalator and the rate of the man in feet per second.
Please show all steps.

Answer by stanbon(48568)

A rushed subway traveler got on an up-escalator and also walked up. After 45 seconds he realized that he had left his briefcase at the bottom. Since he did not have access to a down-escalator, he turned and walked down the up-escalator reaching the bottom in two minutes. If he had traveled up 288 feet, find the rate of the escalator and the rate of the man in feet per second.
Please show all steps.
-----------
Up DATA:
distance: 188 ft ; time=(3/4) minutes ; rate = d/x = 188/(3/4) 250.67 ft/min
--------
Down Data:
distance: 188 ft ; time (5/4)minutes: rate = 188/(5/4) =150.4 ft/min
--------------
Let rate of elevator be e and rate of man be m.
m+e=250.67
m-e=150.4
Add the two equations to get:
2m= 401.07
m=200.33 ft./min (rate of the man)
e=250.67-200.33=50.33 ft/min (rate of the elemator)
==========
Cheers,
Stan H.
m=200.53 ft/min